# Towards Regulatable AI Systems: Technical Gaps and Policy Opportunities

### v1

@@ Abstract
There is increasing attention being given to how to regulate AI systems. As governing bodies grapple with what values to encapsulate into regulation, we consider the technical half of the question: To what extent can AI experts vet an AI system for adherence to regulatory requirements? We investigate this question through two public sector procurement checklists, identifying what we can do now, what we should be able to do with technical innovation in AI, and what requirements necessitate a more interdisciplinary approach.
@@ 1 Introduction
As AI systems become more advanced and integrated into our lives, there has been a corresponding urgency to ensure they align with social values and norms. Legal and regulatory authorities around the world are racing to produce AI regulations [142, 44, 36]. However, the increasing size, generality, opaqueness, and closed nature of present-day AI systems pose significant challenges in achieving this alignment [150, 8]. Even when requirements can be precisely articulated—already a difficult task—there is still the question of whether and how it is possible to check whether an AI system adheres to those requirements.
@@
We consider the following questions: What innovations in AI systems are needed for them to be effectively regulated? And in what areas will innovations in AI methods alone be insufficient, and more interdisciplinary approaches required?
While there are many regulations involving AI, we consider public sector procurement checklists because they are instrumental in shaping societal norms, influence private actors to emulate similar practices, and are relatively explicit. These checklists define the criteria and process for a public sector agency to purchase products or services, including technical criteria that an AI system must satisfy. The technical criteria about AI systems contained in these lists are also present in other regulatory efforts. Thus, improving our ability to vet an AI system against these checklists will improve our ability to regulate AI systems more generally.
Specifically, we closely examine the technical criteria from two existing procurement checklists: the World Economic Forum’s AI Procurement in a Box (WEF) [144] and the Canadian Directive in Automated Decision-Making (CDADM) [56]. The WEF checklist serves as a practical guidebook to unlock the public-sector adoption of AI and has been piloted in various countries, including the UK, Bahrain, the UAE, India, and Brazil [19]. The CDADM is one of the earliest regulations specifically targeting AI systems to be put into operation. It came into effect in April 2019, and requires full compliance by October 2023 for all new systems and by April 2024 for all existing systems. Thus, the question of whether it is possible to build AI systems that meet such checklist criteria is timely.
@@
In the remainder of this document, we first group the technical criteria contained in these two checklists into categories that will be familiar to AI researchers and engineers: (pre-training) data checks, (post-hoc) system monitoring, global explanation, local explanation, objective design, privacy, and human + AI systems. For each category, we briefly summarize existing technical approaches that could be used to construct AI systems that meet those criteria. Next, we identify areas where relevant technical approaches may exist, but additional technical innovation is needed to be able to vet increasingly complex AI systems being used in increasingly varied contexts. For example, the proliferation of large language models comes with a significant difficulty in evaluating them, due to issues including open-endedness and data leakage. While innovative approaches like Holistic Evaluation of Language Models (HELM) [79] and Elo ratings [149] are proposed, the evaluation of language models remains an open question and further technical innovation is needed for effective regulation and oversight. Finally, we briefly outline aspects of these criteria that may seem technical but actually require interdisciplinary approaches to vet.
Throughout this exercise, we assume no concerns about expertise; that is, we assume that there are sufficiently qualified AI and domain experts to review whether the AI system meets the checklist criteria. Our concern is to identify to what extent experts can currently vet AI systems against these regulatory criteria. In doing so, we hope to highlight concrete areas where AI innovation would improve our ability to create regulatable AI systems. There are many subfields in AI. Moreover, AI systems are rapidly advancing, and the kinds of contexts in which they are being used are rapidly growing. Thus, while our list is certainly not comprehensive, we hope it serves as a starting point for AI researchers interested in creating regulatable AI systems. Additionally, this document informs both policy makers and AI engineers on issues where more holistic, interdisciplinary efforts (rather than AI methods alone) are necessary.
@@ 2.3 Areas that require interdisciplinary engagement
The specific metrics that would enable meaningful inference about the quality of the data will depend on the application. Questions around bias and fairness are also inherently multi-faceted and will depend on the use-case. Determining the proper form of consent (opt-in vs. opt-out) is a legal decision. Involvement from social scientists is necessary to assess how different data collection processes include or exclude certain populations. Privacy tensions—what data is retained, what statistics are made public, what kind of access is granted to trusted auditors—must also be resolved within the broader socio-technical context.
@@ 3.3 Areas that require interdisciplinary engagement
At a high level, there will always need to be some kind of decision made about what needs to be monitored or prioritized in a given setting. There will also need to be decisions made about what kinds of safety promises or guarantees are needed e.g. how much shift is considered safe and acceptable, and how much is not. It is crucial to translate the monitored metrics into meaningful implications that enable people to make informed decisions within the broader socio-technical system. For instance, in autonomous driving, comparing monitored metrics against human performance can inform decisions regarding human intervention. Finally, the task of contingency planning for back-ups when models express unexpected or unwanted behaviors also requires an understanding of the broader socio-technical system.
@@ 4.3 Areas that require interdisciplinary engagement
There is a question of what to offer and to whom. For example, releasing the code and environment may allow some people to directly answer their questions. Providing an explanation broadens who can inspect the model, including users and domain experts; however, what information to release, how it should be extracted, and how often during the life cycle of the model that information should be updated will depend on the use context. We will also need mechanisms for people to request more information about a model as new concerns become apparent. Finally, all information release must be balanced with concerns about privacy and trade secrets.
@@ 5.3 Areas that require interdisciplinary engagement
Relatedly, the purpose of the information provided for recourse will vary across contexts. For one task, it may be enough to provide only one recourse, while for others it may be necessary to provide multiple options. In other contexts, the user might benefit from an interactive system to explore different options. For example, they could themselves wish to navigate changes and see if they would result in a favorable loan decision.
Finally, it may be that a recourse generated from a local explanation may not be the appropriate way to assist a user unhappy with a decision. For example, suppose someone is convinced that a voice-based covid test is in error about their disease status. Rather than providing an explanation of the voice features used to make the decision, the appropriate recourse may be to allow that person to take a traditional covid test instead. We also note that certain situations may require a justification (rationale for why a decision is right with respect to laws, norms, and other aspects of the context) rather than explanation (what features the AI used to generate the output).
@@ 6.2 Directions requiring additional AI innovation
  • Robustness to a Variety of Objectives. In some subfields of AI, there is literature on creating agents that perform well across a range of reward functions [91, 101]. This ensures resilience in the face of imperfections in the objective. However, more work is needed to make this process efficient, e.g., for large pre-trained models.
@@
  • Understanding Connections between Objectives and Learnt Model Behavior. Can we efficiently explain how changes in a technical formulation of an objective affect the model behavior? We understand this for certain simple models but not sufficiently for more complex models: e.g., how will changing certain weights in the reward change an autonomous vehicle’s driving? Conversely, can we explain policies in terms of compatible reward functions? Can we efficiently identify where two reward functions may result in different policies in human-understandable terms? Some prior works tries to answer this [48]; however, more analyses will facilitate a more fine-grained design of the reward function to better align with intended objectives.
@@ 6.3 Areas that require interdisciplinary engagement
Creating goals at a policy level requires considering factors such as contextual relevance, attainability, and alignment with overarching desiderata [66]. Ethical concerns associated with the power and impact of AI systems may also be taken into account. Moreover, sometimes even at the policy level, the objective remains unclear, making it more difficult to design proper objectives for the AI systems much less validate and explain them. This issue is evident in the application of AI in criminal justice, where a lack of clear policy goals is common.
@@ 8.3 Areas that require interdisciplinary engagement
Shared human+AI decision-making is an interdisciplinary area involving social science, psychology, cognitive science, etc., [34]. Fortunately, researchers in HCI already have connections to these fields. Furthermore, the design adoption of new tools into workplaces is well-studied in design, human factors research, and management and operations science—and require interdisciplinary teams with appropriate expertise. These interdisciplinary efforts will help inform decisions about whether, how, and which humans to include in the loop, as well as how a system that is expecting human input should respond to inappropriate, slow, or absent input from the human.

### v2

@@ Abstract
There is increasing attention being given to how to regulate AI systems. As governing bodies grapple with what values to encapsulate into regulation, we consider the technical half of the question: To what extent can AI experts vet an AI system for adherence to regulatory requirements? We investigate this question through the lens of two public sector procurement checklists, identifying what we can do now, what should be possible with technical innovation, and what requirements need a more interdisciplinary approach.
@@ 1 Introduction
Regulations represent a fundamental mechanism through which governments fulfill their duty to protect citizens from a variety of threats including health risks, financial fraud, and discriminatory practices. As AI systems become more advanced and integrated into our lives, there has been a corresponding urgency to ensure they align with social values and norms, and that their benefits significantly outweigh any potential harm. In response to this imperative, legal and regulatory bodies globally are engaged in a concerted effort to develop comprehensive AI regulations [177, 58, 47].
However, the increasing size, generality, opaqueness, and closed nature of present-day AI systems pose significant challenges to effective regulation [186, 12]. Even when requirements can be articulated, it remains uncertain if and how we can verify an AI system’s adherence to these standards. A requirement that cannot be checked will not effectively provide protection. If we believe that AI systems should be regulated, then AI systems must be designed to be regulatable.
@@
In this article, we consider the following questions: What innovations in AI systems are needed for them to be effectively regulated? And in what areas will innovations in AI methods alone be insufficient, and more interdisciplinary approaches required?
We explore the answer through the lens of public sector AI procurement checklists, which offer a pragmatic perspective on the broader challenges of regulatable AI systems. These checklists do more than outline criteria for government procurement; they represent concerted efforts to codify regulatory requirements for the adoption of AI systems in the public sector. This positions them uniquely at the intersection of policy-making and practical implementation. The desiderata distilled in these checklists are comprehensive and are reflected in still nascent efforts to regulate AI in the private sector [58, 22]. Furthermore, public sector procurement checklists are among the more developed AI regulations, some having gone through several rounds of refinement [175]. We emphasize that we focus on public sector procurement checklists to make our discussion concrete; the technical innovations needed to satisfy those checklists are relevant to the wider discourse of creating regulatable AI systems.
@@
Specifically, we closely examine the technical criteria from two existing procurement checklists: the World Economic Forum’s AI Procurement in a Box (WEF) [179] and the Canadian Directive on Automated Decision-Making (CDADM) [71]. As illustrated in Figure 1, we first group the technical criteria contained in these two checklists into categories that will be familiar to AI researchers and engineers: (pre-training) data checks, (post-hoc) system monitoring, global explanation, local explanation, objective design, privacy, and human + AI systems. For each category, we briefly summarize existing technical approaches that could be used to construct AI systems that meet those criteria. Next, we identify areas where relevant technical approaches may exist, but additional technical innovation is needed to be able to vet increasingly complex AI systems being used in increasingly varied contexts. For example, the proliferation of large language models comes with a significant difficulty in evaluation, due to factors such as their open-ended nature and data leakage. While innovative approaches like Holistic Evaluation of Language Models (HELM) [103] and Elo ratings [185] are proposed, accurately evaluating these models continues to be an unresolved issue that requires further technical advancements for effective regulation and oversight. Finally, we briefly outline aspects of these criteria that may seem technical but actually require interdisciplinary approaches to vet.
Throughout this exercise, we assume no concerns about expertise; that is, there are sufficiently qualified AI and domain experts to review whether the AI system meets the checklist criteria. Our concern is to identify to what extent experts can currently vet AI systems against these criteria, and provide a (non-comprehensive, but concrete) list of directions for technical innovation to bridge the gap towards regulatable AI systems. If AI systems can be made verified against these checklists, then we will have made significant progress towards creating regulatable AI systems in general.
2 Public Sector AI Procurement Checklists
Figure 1: Connections between the main sections in this article and the high-level CDADM requirements and WEF guidelines. We exclude the WEF guidelines on strategy alignment, multidisciplinary teams, and fair business competition which do not directly motivate CS research.
The public sector uses procurement checklists to ensure its purchased products align with its organizational needs and values. In the context of AI procurement checklists, most pertinent to us are the items related to technical criteria, such as ensuring data quality, privacy, fairness, and appropriate monitoring and oversight. These technical criteria are also contained in broader but more nascent regulatory efforts. [58, 22].
A reasonable question may be why focus on these particular checklists—after all, not all regulations are of the same quality. We note both checklists have gone through extensive review. The directive on automated decision-making marks the Canadian government’s first step toward managing the risks of AI in public administration. Initiated in 2016, this comprehensive effort involved a white paper, workshops, consultation sessions, and the formulation of working groups, drawing upon the expertise of scholars, civil society advocates, and governmental officials [34]. Officially enacted in April 2019, this directive has undergone three rounds of review [24] and was amended in 2021 and 2023 in response to these reviews. It established the groundwork for the broader 2022 Artificial Intelligence and Data Act (AIDA) [32] and a recent guide on the use of generative AI [33].
The ”AI Procurement in a Box” by the World Economic Forum is a set of guidelines, with detailed examples, to assist governments in the creation of public sector procurement regulation. It is the result of a 2018 ”Unlocking Public Sector AI” initiative which included over 200 stakeholders across government, academia, and industry; the efficacy and applicability of the guidelines have been validated through two pilot studies conducted in the UK [180] and Brazil [174].
As noted above, both of the checklists we consider have gone through extensive review. We also spoke to government officials who were using these checklists and carefully examined them with a red-teaming mindset. While no regulatory effort is perfect, we found that these checklists are fairly comprehensive, with perhaps the only technical area lacking being insufficient attention to HCI elements of how the AI would be integrated into its intended environment. We found that their included technical criteria are relevant for broader efforts to create more regulatable AI systems.
For this paper, we aligned and grouped parts of the two checklists with existing AI research topics (outlined in Figure 1). This process required combining information from different checklist sections, highlighting the differences in how AI researchers and policymakers approach the same problems. For example, fairness is a large AI research topic, but interwoven throughout many points in CDADM and WEF, so it does not have its own section here. This paper discusses all major sections of the CDADM but excludes some elements of security and non-expert training in the WEF due to space constraints.
@@ 3.3 Areas that require interdisciplinary engagement
The specific metrics that would enable meaningful inference about the quality of the data will depend on the application [78, 96]. Questions around bias and fairness are also inherently multi-faceted and will depend on the use-case [15]. Determining how data collection for AI systems respects copyright, obtains proper consent (opt-in versus opt-out), and avoids misrepresentation or detriment to the owner’s benefits necessitates input from disciplines such as law, policy, and social science [95, 49, 88, 76]. Privacy tensions—what data is retained, what statistics are made public, what kind of access is granted to trusted auditors—must also be resolved within the broader socio-technical context [50].
@@ 4.3 Areas that require interdisciplinary engagement
At a high level, there will always need to be some kind of decision made about what needs to be monitored or prioritized in a given setting. There will also need to be decisions made about what kinds of safety promises or guarantees are needed e.g. how much shift is considered safe and acceptable, and how much is not, for example in healthcare [60]. It is crucial to translate the monitored metrics into meaningful implications that enable people to make informed decisions within the broader socio-technical system [122]. For instance, in autonomous driving, comparing monitored metrics against human performance can inform decisions regarding human intervention. Finally, the task of contingency planning for back-ups when models express unexpected or unwanted behaviors also requires an understanding of the broader socio-technical system.
@@ 5.3 Areas that require interdisciplinary engagement
There is a question of what to offer and to whom. For example, releasing the code and environment may allow some people to directly answer their questions [127]. Providing an explanation broadens who can inspect the model, including users and domain experts; however, what information to release, how it should be extracted, and how often during the life cycle of the model that information should be updated will depend on the use context. We will also need mechanisms for people to request more information about a model as new concerns become apparent [84]. Finally, all information release must be balanced with concerns about privacy and trade secrets.
@@ 6.3 Areas that require interdisciplinary engagement
Relatedly, the purpose of the information provided for recourse will vary across contexts. For one task, it may be enough to provide only one recourse, while for others it may be necessary to provide multiple options [163, 154]. In other contexts, the user might benefit from an interactive system to explore different options. For example, they could themselves wish to navigate changes and see if they would result in a favorable loan decision.
Finally, it may be that a recourse generated from a local explanation may not be the appropriate way to assist a user unhappy with a decision [6, 3]. For example, suppose someone is convinced that a voice-based covid test is in error about their disease status. Rather than providing an explanation of the voice features used to make the decision, the appropriate recourse may be to allow that person to take a traditional covid test instead. We also note that certain situations may require a justification (rationale for why a decision is right with respect to laws, norms, and other aspects of the context) rather than explanation (what features the AI used to generate the output).
@@ 7.2 Directions requiring additional AI innovation
  • Robustness to a Variety of Objectives. Further research is necessary to create agents that excel across a range of objectives. In RL, this research can strengthen the robustness of learned policies when objectives are not perfectly specified [116, 128]. It also applies to language models, which are trained with the next token prediction task but asked to perform agentic tasks with various constraints and objectives [7], and in different worlds [101].
@@
  • Understanding Connections between Objectives and Learnt Model Behavior. Can we efficiently explain how changes in the objective function impact model behavior? Conversely, can we explain policies in terms of compatible reward functions? Can we efficiently identify where two reward functions may result in different policies in human-understandable terms? Some prior works try to answer this [63]; however, more analysis will help refine the reward function to better match the intended objectives. One further question is disentangling which model behavior is the result of an objective and which is the result of training data. For example, the mix of possibly conflicting beliefs in a text corpus will influence how language models trained on it behave, though all have the same objective. [7]
@@ 7.3 Areas that require interdisciplinary engagement
Creating goals at a policy level requires considering factors such as contextual relevance, attainability, and alignment with overarching desiderata [83]. Ethical concerns associated with the power and impact of AI systems may also be taken into account. Moreover, sometimes the objective remains unclear even at the policy level, making it more difficult to design proper objectives for the AI systems much less validate and explain them. For example, while AI systems are commonly utilized in criminal justice [187, 93], there is often a lack of clarity regarding how they define and measure crime [87], and the data may not accurately reflect the true objectives of judges [44, 137].
@@ 9.2 Directions requiring additional AI innovation
  • Human+AI for Test-time Validation of Large Surface Models. Models with large output surfaces (e.g. LLMs) will be difficult to evaluate via prospective metrics; we need methods for people to assist in their validation at the task-time (e.g. as proposed in [51]).
@@ 9.3 Areas that require interdisciplinary engagement
Shared human+AI decision-making is an interdisciplinary area involving social science, psychology, cognitive science, etc., [43]. Fortunately, HCI research already has connections to these fields [46, 157, 159]. Furthermore, the design adoption of new tools into workplaces is well-studied in design, human factors research, and management and operations science [115, 4, 168]—and require interdisciplinary teams with appropriate expertise. These interdisciplinary efforts will help inform decisions about whether, how, and which humans to include in the loop, as well as how a system that is expecting human input should respond to inappropriate, slow, or absent input from the human.